OpenStudy (anonymous):

evaluate lim sin(x-1)/x^2 +x-2....X>>>1

OpenStudy (raden):

because it be 0/0, so use differential for numerator and denominator it means, limit (x->1) cos(x-1)/(2x+1). now just plug x=1

OpenStudy (anonymous):

everything with x will be 1 right?

OpenStudy (raden):

sorry, i was go just new... yes, it will be cos(1-1)/(2*1+1) = cos0/(2+1) = ...